OpenStudy (anonymous):

please help solve trig equation 2cosxcos2x+sinxsin2x+cosx=0 I need to break it down using identities: sin(2x)=2sinxcosx And ONE of these cos(2x) =cos^2(x)-sin^2(x) cos(2x)=1-2sin^2(x) cos(2x)=2cos^2(x)-1

OpenStudy (campbell_st):

ok... so you can make some substitutions and you get\[2\cos(x)(\cos^2(x) - \sin^2(x)) + \sin(x)(2\sin(x)\cos(x) + \cos(x) =0\] so distribute \[2\cos^3(x)- 2\cos(x)\sin^2(x) + 2\sin^2(x)\cos(x) + \cos(x) = 0\] collect like terms and \[2\cos^3(x) + \cos(x) = 0\] so factoring and you get \[\cos(x)(2\cos^2(x) + 1) = 0\] now I've done the hard work, I'll leave you so solve for x. Hope it helps